# FPT approximations for Capacitated Sum of Radii and Diameters

### v1

@@ 4.1 Node CapSoR
In this section, we prove Theorem 3 for node capacities. Let $\mathcal{I}=((P,\delta),k,\{U_{p}\}_{p\in P})$ be a given instance of non-uniform node CapSoR. Let $\mathcal{C}=\{C_{1},\cdots,C_{k}\}$ be an optimal clustering of $\mathcal{I}$ such that $C_{i}$ has radius $r^{*}_{i}$ with center $o_{i}$ and $r^{*}_{1}\leq\cdots\leq r^{*}_{k}$. Let $O=\{o_{1},\cdots,o_{k}\}$, and let $\text{OPT}=r^{*}_{1}+\cdots+r^{*}_{k}$ be the cost of clustering $\mathcal{C}$. Let $r_{1}\leq\cdots\leq r_{k}$ be an $\varepsilon$-approximation of the optimal radii of $\mathcal{C}$. Algorithm 1 describes the pseudo-code of our algorithm for Theorem 3 that is based on the ideas presented in Section 2. Algorithm 1 makes guesses, and hence, we make the following assumption for brevity. Later, when we proof Theorem 3 in Section 4.1.2, we show how to remove this assumption.
@@
Given a ball $B:={\sf ball}(y,r)$, for $y\in P$ and radius $r$, we say that a cluster $C_{i}\in\mathcal{C}$ is reachable from $B$ if $\delta(y,p)\leq r+2r_{i}$, for all $p\in C_{i}$. Let reachable$(B)$ denote the set of all clusters in $\mathcal{C}$ reachable from $B$. Furthermore, by breaking the ties arbitrary, the cluster with smallest radius in $\emph{{{reachable}}}(B)$ is called anchor for $B$.
@@ 2 $\mathcal{S}\leftarrow(\emptyset,\emptyset)$;
3 Guess $R=(r_{1},\cdots,r_{k})$, an $\varepsilon$-approximation of optimal radii such that $r_{i}\leq r_{i+1}$ for $11\leq i\leq k-$ 1;
@@ 4 Color each cluster $C_{i}\in\mathcal{C}$ red;
7       $C_{i}\leftarrow$ smallest radius red cluster;
@@ 8       SETTLE-CLUSTER $(C_{i})$;
[…]
@@ 15      Delete $y_{t}$ from $F$;
2Let $D_{i}={\sf ball}_{Q}(y_{i},r_{i})$ be the ball returned by FIND-GOOD-DENSE-BALL$(r_{i},(Q,F))$ ;
3 Let $C_{j}$ be an anchor for $D_{i}$;
4 if $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\alpha r_{i}))$ then
@@ 7       Add $(y_{i},(\alpha+2)r_{i})$ to $\mathcal{S}$;
6      Let $C_{j}$ be a cluster with smallest radii in reachable$(D_{i})$;
7       if $C_{i}\notin{\textsf{reachable}}({\sf ball}(y_{i},r_{i}+2r_{j}))$ then  $P^{\prime}\leftarrow P^{\prime}\setminus{\sf ball}(y_{i},r_{i}+2r_{j}))$;
@@ 6      Let $C_{j}$ be a cluster with smallest radii in reachable$(D_{i})$;
9             if $y_{i}$ is not available for $C_{i}$ then  $F^{\prime}\leftarrow F^{\prime}\setminus\{y_{i}\}$;
@@
15fail;
Algorithm 3 FIND-GOOD-DENSE-BALL$(r_{i},(Q,F))$
@@ 4 Add $(y_{i},r_{i})$ to $\mathcal{S}$ and delete $y_{i}$ from $F$;
5
6foreach $C_{t}\in\emph{{reachable}}(D_{i})$ do
7       Add $r_{i}$ to Partitions-of$(C_{t})$;
@@ 4.1.1 Analysis
From Lemma 8, we have that FIND-GOOD-DENSE-BALL$(r_{i},(P,F))$ never fails for every red cluster $C_{i}$ processed by the while loop of Algorithm 1. If $C_{i}$ is a Type-$1$ cluster, then $C_{i}$ is colored black in Step 2 of SETTLE-CLUSTER $(C_{i})$. Otherwise, $C_{i}$ is Type-$2$ cluster and hence it is colored black in Step 4 of EXCHANGE$(D_{i})$. ∎
@@ 4.1.2 Proof of Theorem 3
    Guesses in FIND-GOOD-DENSE-BALL. Each time FIND-GOOD-DENSE-BALL is invoked, the for loop runs $2k$ iterations and each iteration makes two guesses. First one is whether or not $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},r_{i}+2r_{j}))$. For this, we create two branches – one for each outcome. Second guess is whether or not $y_{i}$ is available for $C_{i}$. Again, we create two branches for this guess. Since these two guesses are mutually exclusive, we create at most two branches in each iteration. Hence, we create $2^{2k}$ many branches of execution for each invocation of FIND-GOOD-DENSE-BALL.
@@
    Guesses in SETTLE-CLUSTER. The first guess is of the radius of anchor of $D_{i}$ returned by FIND-GOOD-DENSE-BALL. Hence, we create $k$ branches for this step. Next guess is whether or not $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\alpha r_{i}))$, which results in $2$ branches. Thus, totally, we create $2k$ branches for each invocation of SETTLE-CLUSTER.
@@
Since Algorithm 1 calls SETTLE-CLUSTER at most $k$ times and each execution of SETTLE-CLUSTER creates $2^{O(k)}$ branches of execution, we have that the total running time of Algorithm 1 is bounded by $2^{O(k^{2}+k\log(k/\varepsilon))}n^{O(1)}$.
@@ 1 Let $C^{\prime}_{i}\leftarrow\emptyset$ for $i\in[k]$;
2 Guess $D=(d_{1},\cdots,d_{k})$, an $\varepsilon$-approximation of optimal diameters such that $d_{i}\leq d_{i+1}$ for $11\leq i\leq k-$ 1;
@@ 3 Color each cluster $C_{i}\in\mathcal{C}$ red;
6       $C_{i}\leftarrow$ smallest radius red cluster;
@@ 7       SETTLE-CLUSTER $(C_{i})$;
[…]
@@
2Let $D_{i}={\sf ball}_{Q}(y_{i},d_{i})$ be the ball returned by FIND-GOOD-DENSE-BALL$(d_{i},(Q,F))$ ;
3 Let $C_{j}$ be an anchor for $D_{i}$ ;
4 if $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\beta d_{i}))$ then
@@ 1 $P^{\prime}\leftarrow Q$;
5      Let $C_{j}$ be a cluster with smallest diameter in reachable$(D_{i})$;
6       if $C_{i}\notin{\textsf{reachable}}({\sf ball}(y_{i},d_{i}+d_{j}))$ then  $P^{\prime}\leftarrow P^{\prime}\setminus{\sf ball}(y_{i},d_{i}+d_{j}))$;
@@ 4 Set $C^{\prime}_{i}={\sf ball}(y_{i},d_{i})$;
5
6foreach $C_{t}\in\emph{{reachable}}(D_{i})$ do
7       Add $d_{i}$ to Partitions-of$(C_{t})$;
@@ 5 Hardness of Approximation
The above theorem implies that there is no $f(k,\varepsilon)n^{g(\varepsilon)}$ time algorithm that finds a $(1+\varepsilon)$-approximate solution, for every $\varepsilon>0$, for uniform CapSoR. Such algorithms are called PAS for Parameterized Approximation Scheme [FKLM20] or FPT Approximation Scheme. Therefore, Theorem 2 implies that, assuming gap-ETH, there is no PAS for Uniform CapSoR.
[…]
Towards proving Theorem 1 and Theorem 2, we show a gap preserving reduction from Max3333-SAT5555 to Uniform CapSoR. An instance of Max3333-SAT5555 is same as that of $3$-SAT, but has an additional regularity property that every variable appears exactly in $5$ clauses.
We setup the required preliminaries in Section 5.1. Then, in Section 5.2, we prove Theorem 2 by showing a reduction from Max3333-SAT5555 to uniform CapSoR. Finally, in Section 5.3, we extend this reduction to prove Theorem 1.
@@ 5.1 Preliminaries
The starting point of our reduction is Max3333-SAT5555, a special case of Max3333-SATB, which is defined below.
Definition 13 (Max3333-SATB).
Given a $3$-SAT formula $\phi$ on $N$ variables such that each variable occurs in constant number of clauses, the Max3333-SATB problem asks the maximum number of clauses that can be satisfied by an assignment to the variables. Furthermore, denote by Max3333-SAT5555 when each variable in $\phi$ occurs in exactly $5$ clauses and no variable occurs more than once in a clause.
It is known that Max3333-SATB is APX-hard, and in particular, Max3333-SAT5555 is also APX-hard.
Theorem 7 (Proposition $2.1.2$ of [Fei98]).
Max3333-SAT5555 is APX-hard (in particular there is no PTAS, unless $P=NP$). Specifically, there is some $\varepsilon>0$ such that it is NP-hard to decide if a given instance $\phi$ of Max3333-SAT5555 is satisfiable, or every assignment satisfies at most $(1-\varepsilon)$ fraction of the clauses.
This is hardness of approximation is based on the following polynomial time transformation from Max3333-SATB to Max3333-SAT5555.
Lemma 12 ([Fei98]).
Given an instance $\psi$ of Max3333-SATB on $N$ variables and $M$ clauses, there is polynomial time algorithm that produces an instance $\phi$ of Max3333-SAT5555 on $N^{\prime}$ variables and $M^{\prime}$ clauses such that:
  1. 1.
    $N^{\prime}=O(M)$ and $M^{\prime}=5N^{\prime}/3=O(M)$
  2. 2.
    Every variable in $\phi$ appears in exactly $5$ clauses and every clause contain exactly $3$ variables.
  3. 3.
    If $\psi$ is satisfiable, then $\phi$ is satisfiable.
  4. 4.
    If every assignment of $\psi$ satisfies at most $(1-\delta)M$ clauses for some $\delta>0$, then every assignment of $\phi$ satisfies at most $(1-\varepsilon)M^{\prime}$ clauses, for some $\varepsilon>0$ depending on $\delta$.
@@
There exists constants $\varepsilon^{\prime},\tau>0$ such that no randomized algorithm when given an instance $\phi$ of Max3333-SATB on $N$ variables and $M$ clauses can distinguish the following cases correctly with probability $2/3$ in time $O(2^{\tau N})$.
  • there exists an assignment for $\phi$ that satisfies all $M$ clauses
  • every assignment satisfies $<(1-\varepsilon^{\prime})M$ clauses in $\phi$.
Combining 1 with Lemma 12, we have the following theorem that will be used in our hardness reductions.
Theorem 8 ((Randomized) Gap Exponential Time Hypothesis (gap-ETH) for Max3333-SAT5555).
Assuming gap-ETH, there exists $\varepsilon^{\prime}>0$ such that no randomized algorithm when given an instance $\phi$ of Max3333-SAT5555 on $N$ variables running in time $2^{o(N)}$ can distinguish correctly with probability $2/3$ if it is satisfiable or every assignment satisfies $<(1-\varepsilon^{\prime})$ fraction of clauses.
5.2 Hardness of FPT-approximation for Uniform CapSoR
In this section, we prove Theorem 2. Recall that for an instance $((P,\delta),k,U)$ of uniform CapSoR, we say that a solution $(S\subseteq P,\sigma:P\rightarrow S)$ is feasible if $|S|\leq k$ and $|\sigma^{-1}(s)|\leq U$, for every $s\in S$. Towards proving Theorem 2, we show the following reduction.

### v2

@@ 5.1 Node CapSoR
In this section, we prove Theorem 3 for node capacities. Let $\mathcal{I}=((P,\delta),k,\{U_{p}\}_{p\in P})$ be a given instance of non-uniform node CapSoR. Let $\mathcal{C}=\{C_{1},\cdots,C_{k}\}$ be an optimal clustering of $\mathcal{I}$ such that $C_{i}$ has radius $r^{*}_{i}$ with center $o_{i}$ and $r^{*}_{1}\leq\cdots\leq r^{*}_{k}$. Let $O=\{o_{1},\cdots,o_{k}\}$, and let $\text{OPT}=r^{*}_{1}+\cdots+r^{*}_{k}$ be the cost of clustering $\mathcal{C}$. Let $r_{1}\leq\cdots\leq r_{k}$ be an $\varepsilon$-approximation of the optimal radii of $\mathcal{C}$ (see Section 4.3.2). Algorithm 1 describes the pseudo-code of our algorithm for Theorem 3 that is based on the ideas presented in Section 2. Algorithm 1 makes guesses (using the command guess), and hence, we make the following assumption for brevity. Later, when we proof Theorem 3 in Section 5.1.2, we show how to remove this assumption.
@@
Given a ball $B:={\sf ball}(y,r)$, for $y\in P$ and radius $r$, we say that a cluster $C_{i}\in\mathcal{C}$ is reachable from $B$ if $\delta(y,p)\leq r+2r_{i}$, for all $p\in C_{i}$. Let reachable(B)(B) denote the set of all clusters in $\mathcal{C}$ reachable from $B$. Furthermore, by breaking the ties arbitrarily, the cluster with smallest radius in $\emph{{{reachable}}}(B)$ is called anchor for $B$.
@@ 2 $\mathcal{S}\leftarrow(\emptyset,\emptyset)$;
3 guess $R=(r_{1},\cdots,r_{k})$, an $\varepsilon$-approximation of optimal radii such that $r_{i}\leq r_{i+1}$ for $1\leq i\leq k-1$, and let $(C_{1},\dots,C_{k})$ be the names of the corresponding clusters in optimal;
@@ 5 Set ${\textsf{Partitions-of}}(C_{t})=\emptyset$ for $C_{t}\in\mathcal{C}$;
7 $C_{i}\leftarrow$ smallest radius red cluster according to R;
@@ 8 SETTLE-CLUSTER $(C_{i})$;
12foreach black cluster $C_{t}$ such that $\emph{{{Partitions-of}}}(C_{t})\neq\emptyset$ do
@@ 15  Delete $y_{t}$ from $F$;
[…]
3 guess if $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\alpha r_{i}))$ then
@@ 6   Add $(y_{i},(\alpha+2)r_{i})$ to $\mathcal{S}$;
 guess the radius $r_{j}$ of the anchor of $D_{i}$ ;
    // see Definition 10
6 guess if $C_{i}\notin{\textsf{reachable}}({\sf ball}(y_{i},r_{i}+2r_{j}))$ then  $P^{\prime}\leftarrow P^{\prime}\setminus{\sf ball}(y_{i},r_{i}+2r_{j}))$;
@@
8    guess if $y_{i}$ is not available for $C_{i}$ then  $F^{\prime}\leftarrow F^{\prime}\setminus\{y_{i}\}$;
@@
14fail;
[…]
@@ 4 Add $(y_{i},r_{i})$ to $\mathcal{S}$ and delete $y_{i}$ from $F$;
5 guess ${\textsf{reachable}}(D_{i})$;
6 foreach $C_{t}\in{\textsf{reachable}}(D_{i})$ do
[…]
@@ 5.1.1 Analysis
[…]
@@ 5.1.2 Proof of Theorem 3
    Guesses in FIND-GOOD-DENSE-BALL. Each time FIND-GOOD-DENSE-BALL is invoked, the for loop runs $2k$ iterations and each iteration makes three guesses. First, it guesses the radius $r_{j}$ of the anchor of ${\sf ball}(y_{i},r_{i}+2r_{j})$, which can be simulated by creating $k$ branches corresponding to the elements of the radius profile $R$. Second guess is whether or not $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},r_{i}+2r_{j}))$. For this, we create two branches—one for each outcome. The third guess is whether or not $y_{i}$ is available for $C_{i}$, which again can be simulated by creating two branches for each outcome of the guess. Hence, we create at most $(4k)^{2k}$ branches of execution for each invocation of FIND-GOOD-DENSE-BALL.
@@
    Guesses in SETTLE-CLUSTER. Since we counted the guesses in FIND-GOOD-DENSE-BALL separately, the only guess in this subroutine is whether or not $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\alpha r_{i}))$, which results in additional $2$ branches, for each invocation of SETTLE-CLUSTER.
@@
Since Algorithm 1 calls SETTLE-CLUSTER at most $k$ times and each execution of SETTLE-CLUSTER creates $k^{O(k)}$ branches of execution, we have that the total running time of Algorithm 1 is bounded by $2^{O(k^{2}\log k+k\log(k/\varepsilon))}n^{O(1)}$.
@@ 1 Let $C^{\prime}_{i}\leftarrow\emptyset$ for $i\in[k]$;
2 guess $D=(d_{1},\cdots,d_{k})$, an $\varepsilon$-approximation of optimal diameters such that $d_{i}\leq d_{i+1}$ for $1\leq i\leq k-1$, and let $(C_{1},\dots,C_{k})$ be the names of the corresponding clusters in optimal;
@@ 4 Set ${\textsf{Partitions-of}}(C_{t})=\emptyset$ for $C_{t}\in\mathcal{C}$;
6 $C_{i}\leftarrow$ smallest radius red cluster according to $D$;
@@ 7 SETTLE-CLUSTER $(C_{i})$;
11foreach black cluster $C_{t}$ such that ${\textsf{Partitions-of}}(C_{t})\neq\emptyset$ do
@@
[…]
3 guess if $C_{i}\in{\textsf{reachable}}({\sf ball}(y_{i},\beta d_{i}))$ then
@@ 1 $P^{\prime}\leftarrow Q$;
5 guess the diameter $d_{j}$ of the smallest diameter cluster ${\textsf{reachable}}(D_{i})$;
6 guess if $C_{i}\notin{\textsf{reachable}}({\sf ball}(y_{i},d_{i}+d_{j}))$ then  $P^{\prime}\leftarrow P^{\prime}\setminus{\sf ball}(y_{i},d_{i}+d_{j}))$;
@@ 4 Set $C^{\prime}_{i}={\sf ball}(y_{i},d_{i})$;
5 guess ${\textsf{reachable}}(D_{i})$;
6 foreach $C_{t}\in{\textsf{reachable}}(D_{i})$ do
[…]
@@ 6 Hardness of Approximation
The above theorem implies that there is no $f(k,\varepsilon)n^{g(\varepsilon)}$ time algorithm that can approximate uniform CapSoR to a factor better than $(1+1/e)$. Such algorithms are called PAS for Parameterized Approximation Scheme [FKLM20] or FPT Approximation Scheme. Therefore, Theorem 2 implies that, assuming gap-ETH, there is no PAS for Uniform CapSoR.
In fact, our reduction can be modified to rule out PTAS for Uniform CapSoR, assuming a weaker hypothesis of $\textsf{P}\neq\textsf{NP}$ (restated for convenience). See 1
Our hardness results are based on a polynomial time reduction from Max $k$-Coverage (Max $k$-Cov), where, given a universe $\mathcal{G}$ of elements, a positive integer $k$, and $k$ collections $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ of subsets of $G$, the task is to find $\{S_{1},\dots,S_{k}\}$ such that $S_{i}\in\mathcal{C}_{i},i\in[k]$ that maximizes the number of elements covered. It is known that Max $k$-Cov is NP-hard to approximate to a factor better than (11/e)(1-1/e), even with a promise that there exists $k$ sets covering $\mathcal{G}$ and when every set in the collection has precisely $|\mathcal{G}|/k$ elements.
On a high level, given an instance of Max $k$-Cov, we create the set-element incidence graph of the instance. Additionally, for every collection $\mathcal{C}_{i},i\in[k]$, we form a clique on the vertices corresponding to the sets in $\mathcal{C}_{i}$. Let $G=(V,E)$ be the resulting graph. The points of our CapSoR instance corresponds to the vertices in $V$, and the distance metric is given by the shortest path distance in $G$. Finally, we set the uniform capacity $U=|\mathcal{G}|/k+m$. It is easy to see that when there are $k$ sets $\{S_{1},\dots,S_{k}\}$ such that $S_{i}\in\mathcal{C}_{i},i\in[k]$ that cover $\mathcal{G}$, then we can find a solution for CapSoR with cost $k$, by selecting the points corresponding to the sets $\{S_{1},\dots,S_{k}\}$ as the centers. For the No case, observe that any solution to CapSoR must have clusters of size precisely $U$ due to our construction. This simple observation forces the centers to belong to different collections, yielding the result.
@@ 6.1 Preliminaries
Gap-ETH
@@
There exists constants $\varepsilon^{\prime},\tau>0$ such that no randomized algorithm when given an instance $\phi$ of $3$-SAT on $N$ variables and $O(N)$ clauses can distinguish the following cases correctly with probability $2/3$ in time $O(2^{\tau N})$.
  • there exists an assignment for $\phi$ that satisfies all the clauses
  • every assignment satisfies $(1-\varepsilon^{\prime})$ fraction of clauses in $\phi$.
Max $k$-Coverage
The starting point of our reduction is Max $k$-Cov, where given a universe $\mathcal{G}$, a collection $\mathcal{S}$ of subsets of $\mathcal{G}$, and a positive integer $k$, we want to find $k$ subsets from $\mathcal{S}$ that maximizes the number of covered elements of $\mathcal{G}$. For an element $g\in\mathcal{G}$ and a set $S\subseteq\mathcal{G}$, we say $g$ is covered by $S$, if $g\in S$. Similarly, for subsets $S,T\subseteq\mathcal{G}$, we say $S$ covers the elements of $T$ if $T\subseteq S$. For our reduction, we need a more general setting.
Definition 13 (Max $k$-Cov).
Given a universe $\mathcal{G}$, positive integer $k$, and $k$ collections $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ of subsets of $\mathcal{G}$, the Max $k$-Cov problem asks to find $k$ sets $\{S_{1},\dots,S_{k}\}$ such that $S_{i}\in\mathcal{C},i\in[k]$ that maximizes the number of covered elements of $\mathcal{G}$.
It is known that it is NP-hard to approximate Max $k$-Cov to factors better than (11/e)(1-1/e).
Theorem 7 ([Fei98]).
For every constant $\varepsilon>0$, it is NP-hard to decide whether a given instance of Max $k$-Cov has a feasible solution that covers all the elements of $\mathcal{G}$ or every feasible solution covers at most $(1-1/e+\varepsilon)$ fraction of elements of $\mathcal{G}$. Furthermore, the hardness result holds even when every set in $\mathcal{C}_{i},i\in[k]$ has exactly $|\mathcal{G}|/k$ elements.
In fact, assuming gap-ETH, Manurangsi[ [Man20]], showed the following hardness of approximation for FPT algorithms.
Theorem 8 ([Man20]).
For every constant $\varepsilon>0$ and any function $f$, there is no $f(k)|\mathcal{G}|^{o(k)}$ time algorithm that can decide whether a given instance of Max $k$-Cov has a feasible solution that covers all the elements of $\mathcal{G}$ or every feasible solution covers at most $(1-1/e+\varepsilon)$ fraction of elements of $\mathcal{G}$, unless gap-ETH  fails. Furthermore, the hardness result holds even when every set in $\mathcal{C}_{i},i\in[k]$ has exactly $|\mathcal{G}|/k$ elements.
6.2 Reduction from Max $k$-Cov to Uniform CapSoR
In this section, we show the following result.
Theorem 9.
Given any constant $\varepsilon>0$ and an instance $\mathcal{I}=((\mathcal{G},\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}),k)$ of Max $k$-Cov with $|\mathcal{C}_{i}|=m,i\in[k]$ such that $m\leq\nicefrac{{n^{\prime}}}{{2k^{2}}}$, where $n^{\prime}=|\mathcal{G}|$ and every set in $\mathcal{C}_{i},i\in[k]$ has exactly $\nicefrac{{n^{\prime}}}{{k}}$ elements, one can compute in ${\rm poly}(n^{\prime})$ time an instance $\mathcal{J}=((P,\delta),k,U)$ of uniform CapSoR with $n:=|P|=n^{\prime}+mk$ points and uniform capacity constrain $U=n/k$, such that
  • If $\exists\mathcal{T}=\{T_{1},\dots,T_{k}\mid T_{i}\in\mathcal{C}_{i},\forall i\in[k]\}$ such that $\mathcal{T}$ covers $\mathcal{G}$, then there exist a feasible solution $(S,\sigma)$ with ${\rm cost}(S,\sigma)\leq k$
  • If $\forall\mathcal{T}=\{T_{1},\dots,T_{k}\mid T_{i}\in\mathcal{C}_{i},\forall i\in[k]\}$, $\mathcal{T}$ covers strictly less than $(1-1/e+\varepsilon)n^{\prime}$ elements of $\mathcal{G}$, then then for every feasible solution $(S,\sigma)$ it holds that ${\rm cost}(S,\sigma)>(1+1/e-\varepsilon)k$.
Proof.
For every element and every set in $\mathcal{I}$, create a point in $P$. We call points corresponding to elements, as element points, and points corresponding to sets, as set points. Formally, for every element $e\in\mathcal{G}$, create an element point $p_{e}$ in $P$, and for every set $T\in\bigcup_{i\in[k]}\mathcal{C}_{i}$, create a set point $p_{T}$ in $P$. Note that $n=|P|=n^{\prime}+mk$. Define the distance function $\delta$ as the shortest path metric of an unweighted graph $G=(V,E)$, where $V=P$, and $E$ has the following edges. $(i)$ For every $\mathcal{C}_{i},i\in[k]$, there is an edge between every pair of set points $p_{T},p_{T^{\prime}}\in P$ such that the corresponding sets $T,T^{\prime}\in\mathcal{C}_{i}$. $(ii)$ For every set point $p_{T}\in P$, corresponding to set $T$, there are edges between $p_{T}$ and the element points corresponding to the elements contained in $T$. Finally, set $U=n/k=n^{\prime}/k+m$.